# RETTA: Retrieval-Enhanced Test-Time Adaptation for Zero-Shot Video Captioning

### v2

@@ 4.4 Ablation Study
Table 4: Performance with different our proposed component. Experiments are conducted on the VATEX dataset.
@@
Effect of video retriever. In order to study the impact of video retrievers on our method, we use different configurations of XCLIP as our retriever. The results are present in Table 6. We observe that when XCLIP utilizes a ViT [47] with a size of B/16 as the backbone, it achieves an approximate improvement of 2.2% in the CIDEr score compared to using B/32 (Row 2 v.s. Row 1). The Kinetics-600 [25] dataset is the extended version of Kinetics-400 [48], containing more training data. Therefore, XCLIP pre-trained on Kinetics-600 brings a further improvement of 1.7% compared to XCLIP pre-trained on Kinetics-400 (Row 3 v.s. Row 2). Row 4 achieves the best performance on all metrics. It is trained with 16 frames per video and outperforms Row 3 using 8 frames. Therefore, in the other experiments conducted in this paper, we use the XCLIP in Row 4 as our default retriever.
Effect of the number of retrieved sentences. We study the effect of using different numbers of retrieved sentences. Figure 4 shows the difference in video captioning performance as $K$ varies from 5 to 20. We find that retrieving 15 sentences offers better CIDEr scores and BLEU@4 scores compared to other numbers. It can be concluded that too many or too few retrieval sentences can degrade the performance. This is because too many retrieved sentences may introduce too much noise while too few retrieved sentences cannot provide sufficient information.
@@
Effect of the number of learnable tokens. We study the effect of the number of learnable tokens, and Figure 5 shows that 5 is the better choice than other numbers. In Figure 5(a), the CIDEr score initially decreases as the number increases from 1 to 3, then rises to its highest point from 3 to 5, followed by another decrease from 5 to 7. In Figure 5(b), the Bleu@4 score first increases as the number goes from 1 to 5, then decreases from 5 to 7. Therefore, we use 5 as our default number of soft tokens.
(a) CIDEr under different numbers of retrieved sentences
(b) BLEU@4 under different numbers of retrieved sentences
Figure 4: Performance (CIDEr and BLEU@4 score) comparison under different numbers of retrieved sentences on the VATEX dataset.
(a) CIDEr under different numbers of soft tokens
(b) BLEU@4 under different numbers of soft tokens
Figure 5: Performance (CIDEr and BLEU@4 score) comparison under different numbers of soft tokens on the VATEX dataset.
@@
Effect of the frequency size of the word-granularity loss. In Fig. 6, we conduct experiments on VATEX to explore the impact of the frequency size $L$ of word-granularity loss. From the figure, we observe that the CIDEr score first rises before reaching the moderate number (i.e., $L=5$), and then starts to decrease slightly. The main reason is that when $L$ is large, some noisy words are introduced. On the contrary, if $L$ is small, some essential words may be neglected. In Fig. 6(b), a similar trend is observed in the BLEU metric as well. Therefore, we select an appropriate number (i.e., $L=5$) as the frequency size.
(a) CIDEr under different frequency size
(b) BLEU@4 under different frequency size
Figure 6: Performance (CIDEr and BLEU@4 score) comparison under different frequency sizes on the VATEX dataset.
@@
Effect of the number of candidate words As mentioned in Sec. 3.2, to enhance efficiency, we compute the loss using the top 100 candidate words based on the original probability distribution obtained by the language model. In Figure 7(a) and Figure 7(b), we conduct experiments on VATEX to study the impact of the number of candidate words. From Figure 7(a), we find that using the top 250 candidate tokens achieves the highest CIDEr score compared to other numbers, with the top 100 candidate tokens achieving the second-best CIDEr score. In contrast, in Figure 7(b), we observe that the average time consumed per video increases roughly linearly with the number of candidate tokens. Compared to using 100 tokens, using 250 tokens takes about twice as long. To balance performance and time consumption, we set the number of candidate words to 100.
(a) CIDEr under different numbers of candidate words
(b) Average runtime under different numbers of candidate words
Figure 7: CIDEr and Average runtime per video under different numbers of candidate words on the VATEX dataset.
@@
Effect of hard prompt set In our method, we randomly sample a phrase from a predefined set as the hard prompt. The set we designed includes four phrases: “Video showing”, “Video of”, “Video shows” and “Video describes”. In Table 7, we compare the performance of our method under different sizes of hard prompt sets. Row 1 displays the worst CIDEr score in the absence of any hard prompts. Using “Video showing” as a hard prompt consistently results in a CIDEr improvement of about 1.7% (Row 2 v.s. Row 1). The inclusion of “Video showing” and “Video of” in the hard prompt set leads to a CIDEr increase of around 0.8% (Row 3 v.s. Row 2). With the addition of “Video shows”, the CIDEr score in Row 4 leads by an absolute 0.3 over Row 3. Therefore, in our other experiments, the phrases covered in Row 4 are set as our default hard prompt set.
@@
4.5 Qualitative Results
Several video captioning qualitative results are shown in Fig. 8. We can see that the caption produced by our method is more specific than other methods. For example, the video in Fig. 8 (a) shows two men playing ping pong. Our method accurately describes “The two boys playing ping pong in a table tennis court”. In Fig. 8 (b), although the fully supervised method HMN describes “A man is riding a motorcycle”, our method more accurately identifies the “ocean wave”. In Fig. 8 (c), our method effectively describes “The tutorial instructions for downloading songs”, whereas the compared zero-shot method EPT merely mentions “download” but does not capture “songs”. In Fig. 8 (d), only our method correctly describes the video content “Cooking shrimp in oil”. In contrast, the generated caption “A person is preparing a recipe” by HMN is too general. These results further show the zero-shot ability of our proposed method.
Fig. 9 also shows several zero-shot image captioning qualitative results. In Fig. 9 (a), our method identifies “laptop computer” and correctly describes the image “A baby typing on laptop computer”. In Fig. 9 (b), while three methods all recognize “pizza”, our method additionally captures “plate”. In Fig. 9 (d), our method vividly generates “Elephant slide in the park”.
@@ 5 Conclusion
In this paper, we make the first attempt to explore the combination of a video-text retrieval model, an image-text matching model, and a large language model under the zero-shot and fast adaptation setting for video captioning. Specifically, we use learnable tokens as an information medium connecting these frozen models. With the help of our carefully designed loss functions, these tokens are able to convey video information to the LLM in a way that serves as suitable soft prompts. The soft target-based inference-time token update ensures fast adaptation. Our proposed framework is conceptually simple and easy to implement. The extensive experiments on three benchmarks demonstrate its favorable zero-shot ability. We will make the code publicly available and hope our work can inspire future research in other areas.

### v3

@@ 4.4 Ablation Study
Table 5: Performance with different our proposed component. Experiments are conducted on the VATEX dataset. $\mathcal{L}_{retrieval}^{S}$, $\mathcal{L}_{retrieval}^{W}$, and $\mathcal{S}_{clean}$ represent our proposed losses at the sentence and word levels, as well as the sentence-cleaning strategy, respectively.
@@
Effect of video retriever. In order to study the impact of video retrievers on our method, we use different configurations of XCLIP as our retriever. The results are present in Table 7. We observe that when XCLIP utilizes a ViT with a size of B/16 as the backbone, it achieves an approximate improvement of 2.2% in the CIDEr score compared to using B/32 (Row 2 v.s. Row 1). The Kinetics-600 dataset is the extended version of Kinetics-400, containing more training data. Therefore, XCLIP pre-trained on Kinetics-600 brings a further improvement of 1.7% compared to XCLIP pre-trained on Kinetics-400 (Row 3 v.s. Row 2). Row 4 achieves the best performance on all metrics. It is trained with 16 frames per video and outperforms Row 3 using 8 frames. Therefore, in the other experiments conducted in this paper, we use the XCLIP in Row 4 as our default retriever.
Figure 5: Performance comparison on the VATEX dataset under different settings: (a)–(b) retrieved sentence count, (c)–(d) soft token count, (e)–(f) frequency size, and (g)–(h) candidate word count.
Effect of the number of retrieved sentences. We study the effect of using different numbers of retrieved sentences. Figure 5 (a) and (b) show the difference in video captioning performance as $K$ varies from 5 to 20. We find that retrieving 15 sentences offers better CIDEr scores and BLEU@4 scores compared to other numbers. It can be concluded that too many or too few retrieval sentences can degrade the performance. This is because too many retrieved sentences may introduce too much noise while too few retrieved sentences cannot provide sufficient information. A similar observation was made in [48], which notes that additional retrieved sentences may bring in unrelated descriptions and semantic noise, ultimately harming caption quality. Therefore, setting K=15 results in the best captioning performance.
@@
Effect of the number of learnable tokens. We study the effect of the number of learnable tokens, and Figure 5 (c) and (d) show that 5 is the better choice than other numbers. In Figure 5 (c), the CIDEr score initially decreases as the number increases from 1 to 3, then rises to its highest point from 3 to 5, followed by another decrease from 5 to 7. In Figure 5 (d), the Bleu@4 score first increases as the number goes from 1 to 5, then decreases from 5 to 7. Therefore, we use 5 as our default number of soft tokens. A similar choice was made in the work [19], which uses 4 learnable tokens.
@@
Effect of the frequency size of the word-granularity loss. In Figure 5 (e) and (f), we conduct experiments on VATEX to explore the impact of the frequency size $L$ of word-granularity loss. From the Figure 5 (e), we observe that the CIDEr score first rises before reaching the moderate number (i.e., $L=5$), and then starts to decrease slightly. The main reason is that when $L$ is large, some noisy words are introduced. On the contrary, if $L$ is small, some essential words may be neglected. In Figure 5 (f), a similar trend is observed in the BLEU metric as well. A similar choice was also made in [10] when selecting the number of retrieved entities. Moreover, comparative results show that increasing the number of retrieved entities beyond 5 fails to improve metrics and instead causes performance to drop. Therefore, we select an appropriate number (i.e., $L=5$) as the frequency size.
@@
Effect of the number of candidate words As mentioned in Sec. 3.2, to enhance efficiency, we compute the loss using the top 100 candidate words based on the original probability distribution obtained by the language model. In Figure 5 (g) and Figure 5 (h), we conduct experiments on VATEX to study the impact of the number of candidate words. From Figure 5 (g), we find that using the top 250 candidate tokens achieves the highest CIDEr score compared to other numbers, with the top 100 candidate tokens achieving the second-best CIDEr score. In contrast, in Figure 5 (h), we observe that the average time consumed per video increases roughly linearly with the number of candidate tokens. Compared to using 100 tokens, using 250 tokens takes about twice as long. To balance performance and time consumption, we set the number of candidate words to 100.
@@
Effect of hard prompt set In our method, we randomly sample a phrase from a predefined set as the hard prompt. The set we designed includes three phrases: “Video showing”, “Video of”, and “Video shows”. In Table 8, we compare the performance of our method under different sizes of hard prompt sets. Row 1 displays the worst CIDEr score in the absence of any hard prompts. Using “Video showing” as a hard prompt consistently results in a CIDEr improvement of about 1.7% (Row 2 v.s. Row 1). The inclusion of “Video showing” and “Video of” in the hard prompt set leads to a CIDEr increase of around 0.8% (Row 3 v.s. Row 2). With the addition of “Video shows”, the CIDEr score in Row 4 leads by an absolute 0.3 over Row 3. Therefore, in our other experiments, the phrases covered in Row 4 are set as our default hard prompt set. Inspired by ZeroCap’s use of the manually crafted “Image of” prompt [25], our predefined set of hard prompts was also manually designed and consists of three simple and intuitive phrases. During inference, following prior work [39], the hard prompts are sampled using a uniformly random strategy. Table 9 compares the effect of different sampling strategies, including fixed sampling of each specific prompt and uniformly random sampling. We find that using any of the three individual fixed prompts yielded very similar performance (CIDEr scores of 23.2–23.3), indicating that the model does not show a strong preference for any specific phrase. Thus, weighting among prompts is unnecessary. Moreover, random sampling achieves a mean of 24.0 and a standard deviation of 0.3, reflecting a stable and slight improvement over fixed prompts. Supported by prior studies [39] and our experimental results, this strategy appears to be a reasonable choice.
@@
Table 9: Comparisons of CIDEr performance under different prompt sampling strategies. For random sampling, we report the mean and standard deviation across multiple runs.
Strategy Fix sampling (video showing) Fix sampling (video of) Fix sampling (video shows) Uniformly random sampling
CIDEr 23.3 23.2 23.2 24.0 $\pm$ 0.3
@@
4.5 Qualitative Results
Several video captioning qualitative results are shown in Figure 6. We can see that the caption produced by our method is more specific than other methods. For example, the video in Figure 6 (a) shows two men playing ping pong. Our method accurately describes “The two boys playing ping pong in a table tennis court”. In Figure 6 (b), although the fully supervised method HMN describes “A man is riding a motorcycle”, our method more accurately identifies the “ocean wave”. In Figure 6 (c), our method effectively describes “The tutorial instructions for downloading songs”, whereas the compared zero-shot method EPT merely mentions “download” but does not capture “songs”. In Figure 6 (d), only our method correctly describes the video content “Cooking shrimp in oil”. In contrast, the generated caption “A person is preparing a recipe” by HMN is too general. These results further show the zero-shot ability of our proposed method.
Figure 7 also shows several zero-shot image captioning qualitative results. In Figure 7 (a), our method identifies “laptop computer” and correctly describes the image “A baby typing on laptop computer”. In Figure 7 (b), while three methods all recognize “pizza”, our method additionally captures “plate”. In Figure 7 (d), our method vividly generates “Elephant slide in the park”.
@@ 5 Conclusion
In this paper, we make the first attempt to explore the combination of a video-text retrieval model, an image-text matching model, a text alignment model, and a text generation model under the zero-shot and fast adaptation setting for video captioning. Specifically, we use learnable tokens as an information medium connecting these frozen models. With the help of our carefully designed loss functions, these tokens are able to convey video information to the LLM in a way that serves as suitable soft prompts. The soft target-based inference-time token update ensures fast adaptation. Our proposed framework is conceptually simple and easy to implement. The extensive experiments on three benchmarks demonstrate its favorable zero-shot ability. However, our work still has certain limitations. For example, it currently relies on a content-rich corpus as its database. This dependency may lead to less accurate descriptions generated by the model when the corpus is not rich in content. Therefore, constructing a large-scale and diverse corpus is necessary to further improve the model’s descriptive capability. In the future, we plan to construct a large-scale corpus to further enhance the accuracy of the model in generating visual descriptions. In addition, we aim to extend our method to cross-domain video captioning tasks [50], which may further enhance its generalization ability in broader cross-modal understanding scenarios. We will make the code publicly available and hope our work can inspire future research in other areas.
6 Acknowledgement
This work was partly supported in part by the National Natural Science Foundation of China under Grants 62272438, in part by Beijing Natural Science Foundation L25700, and in part by the Fundamental Research Funds for Central Universities (E2ET1104). Yuankai Qi, Amin Beheshti, and Quan Z. Sheng are not supported by the above-mentioned funds.